Pizzeria Duo has current assets of $6943, fixed assets of $24847, current liabilities of $6279, and long-term debt of $8555. What is the value of Pizzeria Duo's Owners Equity? (Round to nearest penny and do not enter commas, e.g. 1234.56)
Answer #1

Total assets=current assets+Fixed assets

=6943+24847=31790

Total liabilities=current liabilities+long term debt

=(6279+8555)=14834

Total assets=Total liabilities+Total equity

Total equity=31790-14834

=$16956